OpenStudy (anonymous):

Dividing and Multiplying with like bases. If the variables are the same, but the coefficients are not. Does It still count as "Like Bases?"

OpenStudy (anonymous):

no

OpenStudy (anonymous):

The variable bases are separate from numerical factors. Each number or variable is considered separately. So the answer is yes.

OpenStudy (anonymous):

\[2a ^{8}\times a ^{5}\times7a ^{2}\] So I can just add the exponents?

OpenStudy (anonymous):

For example \[12a^5b^3/(6a^2b^7)\] would simplify as \[2a^3/b^4\]The 12 and 6, the a's and b's are each simplified independently of the the others.

OpenStudy (anonymous):

Yes. Multiply the 2*7=14, and add the exponents 8+5+7=20, thus the answer would be \[14a^{20}\]

OpenStudy (anonymous):

You truly are a lifesaver. Thanks :)

OpenStudy (anonymous):

Sorry! I misread the third exponent as the 7 instead of 2. The exponent would be 15.

OpenStudy (anonymous):

Lol thanks just confirming.